40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sun Mar 11 18:43:07 2018, Craig, DailyProgress, PEM, Notes from today's work 
    Reply  Mon Mar 12 01:25:23 2018, Craig, DailyProgress, PEM, Notes from today's work 
       Reply  Mon Mar 12 15:22:42 2018, Craig, Jaime, DailyProgress, PEM, Notes from today's work 
    Reply  Thu Mar 15 10:29:04 2018, awade, DailyProgress, scatter, Addressing 500 Hz scatter pickup 2018-03-14_16.58.21.jpg2018-03-14_20.40.18.jpg2018-03-14_16.58.13.jpg
Message ID: 2131     Entry time: Mon Mar 12 15:22:42 2018     In reply to: 2130
Author: Craig, Jaime 
Type: DailyProgress 
Category: PEM 
Subject: Notes from today's work 

Note on (2) from below:
It seems that fb4's system time is fine.  I have synced up ws1 as well so that both system times are okay by installing ntp: $ sudo apt-get install ntp.  This worked automatically on ws1.
Jaime and I restarted fb4, but only ten minutes after doing so, fb4's frame time was three minutes behind.  Jaime suspects this is because the frame gpstime (C4:DAQ-DC0_GPS) is unregulated.  Jaime is currently working on preventing fb4 frame time drift.

Quote:

More notes:
1) The AOMs temperature drift are probably responsible for the "breathing" we're seeing in the Trans DC PDs.  We should get the ISSs working again as soon as possible.  Probs need thermal hats for the AOMs to stop thermal drift.

2) I am unable to access current CTN Lab data on fb4 using my python method from a few elogs ago.  (Now to get on our workstation you have to access through port 22: $ ssh -Y controls@131.215.115.216 -p 22)
I am also unable to access data using data viewer directly on fb4...  It says

Connecting to NDS Server localhost (TCP port 8088)
Connecting.... done
No data found

read(); err=Bad file descriptor
read(); err=Bad file descriptor
T0=18-03-10-06-36-21; Length=10800 (s)
No data output.

All of this used to work.  This needs to be resolved tomorrow.
Additionally, I think that the fb4 frame times are off by two days, because when I press "Time Now" in Data Viewer it gives me some time from March 10th (It's March 12th right now).  This is especially confusing because the $ date command on fb4 gives the correct time, but the frames are not at the correct time. 

3) I created nine new channels for autolocker tuning in real time.  awade should use these in autolocker.py in Git/cit_ctnlab/ctn_scripts/.  I already put the channel names in the three .ini files, ALConfig_NCAV.ini, ALConfig_SCAV.ini, and ALConfig_NPMC.ini.  Will need to restart acromag to activate these channels.
Channels:
On acromag1, in ~/modbus/db/PMCInterfaceControls.db
C3:PSL-NCAV_PMC_AUTOLOCKER_SLOW_SWEEP_BOTTOM
C3:PSL-NCAV_PMC_AUTOLOCKER_SLOW_SWEEP_TOP
C3:PSL-NCAV_PMC_AUTOLOCKER_SLOW_SWEEP_STEPSIZE
On acromag1, in ~/modbus/db/AutoLockerSoftChannels.db
C3:PSL-NCAV_FSS_AUTOLOCKER_SLOW_SWEEP_BOTTOM
C3:PSL-NCAV_FSS_AUTOLOCKER_SLOW_SWEEP_TOP
C3:PSL-NCAV_FSS_AUTOLOCKER_SLOW_SWEEP_STEPSIZE
C3:PSL-SCAV_FSS_AUTOLOCKER_SLOW_SWEEP_BOTTOM
C3:PSL-SCAV_FSS_AUTOLOCKER_SLOW_SWEEP_TOP
C3:PSL-SCAV_FSS_AUTOLOCKER_SLOW_SWEEP_STEPSIZE

Quote:

Today I buzzed the table and determined there was a strong 500 Hz dirty resonance on the first steering mirror after the PMC. 
This caused me to go around tightening bolts everywhere, including the offending steering mirror and the optics around it.  This did not reduce the resonance.
I tightening the PMC REFL steering mirror as well, and this caused a misalignment onto the PMC REFL PD.  This took me a little while to figure out why the North path refused to lock.  I realigned the PMC REFL steering mirror into the PD.
After I got the North PMC locking again, the North path itself was not locking anymore.  I reranged the autolocker slow volts, but this did not help. 
Turns out the North Trans PD threshold voltage was not high enough.  This is likely because of the bolt tightening, causing some slight misalignment into the North cavity, lowering the overall circulating power in the cavity.  I lowered the autolocker threshold from 1.1 volts to 1.0 volts, and aligned the North Trans PD.  We need to rescan the North cavity to get better alignment/mode matching, but I'm gonna put this off until we replace this offending 500 Hz post-PMC steering mirror.
While I was realigning the Trans PD, I noticed that even touching the trans optics tables causes large ~1Hz oscillations in the trans voltage.  This is definitely exacerbating any scattering problem we have.  Also, the Trans PD output for both paths is "breathing", going up and down with a period of about a minute.  This is bad for our autolocker's threshold.  It's possible that we should build two periscopes for the north and south paths to eliminate these elevated tables which cause coherent oscillations on all trans optics.  We could copy Tara's front periscope design.

 

 

ELOG V3.1.3-